Inhaltsverzeichnis

Christian Morgenstern, Varulven; Christian Morgenstern, The Banshee (An Approach); Christian Morgenstern, Der Werwolf . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Christian Morgenstern, Varulven; Christian Morgenstern, The Banshee (An Approach); Christian Morgenstern, Der Werwolf

2/3

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,

Varulven The Banshee (An Approach)

Der Werwolf

tradukita de Sven Collberg tradukita de Max Knight

En varulv lopp från fru m.m. en vacker afton samt begav sig till en bymagisters grav och bad: “Var snäll och kon-

jugera.

One night, a banshee slunk away from mate and child, and in the

gloom went to a village teacher’s tomb, requesting him: “Inflect me,

pray.”

Ein Werwolf eines Nachts ent-

wich von Weib und Kind und sich

begab an eines Dorfschullehrers Grab und bat ihn: �Bitte, beuge

mich!“

mig!” Byskolläraren stod upp på blecknamnsskyltens

mässingknopp med svar till den, som satt på

pass med tåligt korslagd tass på

tass.

The village teacher climbed up

straight upon his grave stone with its pla-

te and to the apparition said who meekly knelt before the

dead:

Der Dorfschulmeister stieg

hinauf auf seines Blechschilds Mes-

singknauf und sprach zum Wolf, der seine

Pfoten geduldig kreuzte vor dem To-

ten:

“Jag var-ulv 1:a pers. i sing., du, han, hon, den, det var-

ulv. Så vi voro-ulv, I voren. På de voro-ulv kom ingenting.”

“The banSHEE, in the subject’s

place; the banHERS, the possessive ca-

se. The banHER, next, is what they

call objective case–and that is all.”

�Der Werwolf“ – sprach der gu-

te Mann, �des Werwolfs, Genitiv so-

dann, dem Werwolf, Dativ, wie

man’s nennt, den Werwolf, - damit hat’s ein

End!“

Varulven myste gladåt hur pers. hade följts av sing. och

plur. “Fast”, bad han, “kan det inte

ges ens en enda pers. i tempus pre-

sens?”

The banshee marveled at the ca-

ses and writhed with pleasure, ma-

king faces, but said: “You did not add, so far, the plural to the singular!”

Dem Werwolf schmeichelten

die Fälle er rollte seine Augenbälle. �Indessen“, bat er, �füge doch zur Einzahl auch die Mehrzahl

noch!“

…  …  …

Christian Morgenstern, Varulven; Christian Morgenstern, The Banshee (An Approach); Christian Morgenstern, Der Werwolf

3/3

Bekänna måste vår magister, att därvidlag fanns vissa bri-

ster. En varulv sine imperfecto var contradictio in adjecto.

The teacher, though, admitted

then that this was not within his ken. “While bans are frequent”, he ad-

vised, “A she cannot be plurized.”

Der Dorfschullehrer aber mus-

ste gestehn, dass er von ihr nichts

wusste. Zwar Wölfe gab’s in großer

Schar, doch �Wer“ gab’s nur im Sinu-

lar.

Varulven satt med tårad lins. Han visste att han fanns och

finns. Därtill kom också fru m.m. som även måste existera.

The banshee, rising clammily, wailed: “What about my family?” Then, being not a learned crea-

ture, said humbly “Thanks” and left

the teacher.

Der Wolf erhob sich

tränenblind – er hatte ja doch Weib und

Kind! Doch da er keine Gelehrter

eben, so schied er dankend und erge-

ben.

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Svedan de Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Verkinto de tiu ĉi Germana poemo estas Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

http://www.poezio.net/version?poem-id=825&version-id=1654,1660,1652